two cyclists start from the same point and ride in opposite deirections. ONe cyclist rides twice as fast as the the other. In 3 hours they are 81 miles apart. Find the rate of each cyclist.
Answer by checkley77(12569)

3x+3*2x=81
3x+6x=81
9x=81
x=81/9
x=9 mph for the slower biker.
2*9=18 mph for the faster biker.
Proof:
3*9+6*9=81
27+54=81
81=81